RC Update and Run Plan for Wednesday, March 11, 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are M-F at 4:15 PM, 2nd floor Counting House Meeting Room Bluejeans Connection

Tuesday, March 10, 2020.

General updates:

  • Continuing the program of longitudinal target production, nominal HMS/SHMS -3.5/-3.4 GeV data with 3He target.
  • 4 target polarization measurements since Tuesday SWING: 51.1%, 51.0%, 51.9%, and 53.4% . All longitudinal.
  • Let the new RC, Whitney Armstrong, and Walter Kellner know about any non-pivot work you'd like to take place during the March 13-20 down in Hall C. (Please CC Brad as well.) So far on the list we've got:
      • a 3He compressor checkout for the x > 1 experiment next year (Meekins/RadCon);
      • a hopefully straightforward SHMS dipole PS relay replacement followed by testing (Lassiter/Segal);
      • replace a gauge in the SHMS detector hut and put a lead blanket around it (Kenyon);
      • and an important SOLID detector test the scope of which is defined in the HClist http://devweb.acc.jlab.org/CSUEApps/hclist/task/1809 (Malace et al).
  • Whit's got the RC phone. Also: whit@jlab.org .

Current Run Plan:

Wednesday DAY to Thursday DAY

  • ONGOING Resume longitudinal production data taking with target at 180 deg, take NMR every 5 hours.

Thursday DAY to Friday DAY

  • Moeller measurement starting 08:30 on Thursday.
  • 1 hour study by Mike T on the systematics of beam loss versus the C slit setting.
  • Resume longitudinal production data taking with target at 180 deg, take NMR every 5 hours.
  • Hall C one week shutdown for magnet rotation starting Friday 08:00 ish.

Target coil rotation planned for March 13--20

  • Nominal 1-week downtime in Hall C to rotate target coils, install new cell, and transition to d2n program

Notes for the Future

  • At d2n startup, it would be a good idea to quantitatively check the dispersion at the target. I don't think we need a tight spec, but if it's as sloppy as 1 cm/% (hence a decent spectrometer), then we're more at risk for things like how much the beam moves when the locks are off, or scraping from an off-energy tail on the beam.
  • 'Filler' runs for next spin up / target-down period (Don't want to forget about this!):
    • DONE 1 hour run with target out for measure of the background just from the beryllium windows. (Needs SHMS, btw.)

Pending Issues (new stuff in red)

  • A replacement laser interlock jumper needs to be made and installed as soon as possible.
  • Maybe we can ask Mike Fowler for a "cycled" flag on the magnet screens for Xmas?
  • pNMR issues
    • Software/driver issue needs to be resolved
    • Checkout during next non-production down time (Moller or SHMS fix)
  • EPR problems [1] saw some progress on Tuesday [2], but issues remain
    • Tentative plan now is to continue the d2n (and A1n) program by replacing the photo-diode that is placed directly on the pumping chamber (we have many spares)
    • It would be good to write up a procedure to streamline this replacement procedure for us and RadCon
  • Kepco power supply commissioning/installation (Todd, Bill) (Bill got a filter design from the vendor and has ordered parts.)
  • Verify that calibration constants and input values associated with NMR and other polarimetry are consistent and will be correct when the software is restarted [3] [4] [5] [6] [7] (Bill, Junhao)
  • Persistent BLM & ion chamber trips
    • General activation, weird BLM issues (https://logbooks.jlab.org/entry/3792326), Ad-hoc ion chamber threshold adjustments (https://logbooks.jlab.org/entry/3792095)
    • NOTE: Shift crew should NOT give authorization to MCC change ion chamber thresholds. If they ask, tell Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)
    • Beam Studies program will investigate some of these issues (most pressing is significant activation + beam loss in the arcs)
  • Contact with Target readbacks through AnywhereUSB USB bridge still drop out on occasion
    • Initial replacement of the USB bridge failed (did not seem to help underlying issue, then bridge stopped working), replaced with original device
  • OPs wished to reboot hung iochc10 [8], noted in December as well [9]
    • This runs the Moller cryo controls and should not be rebooted without expert oversight (does not impact operations). D. Gaskell is looking into getting this addressed properly with OPs.

Opportunistic Access Jobs

None at the moment.


Standing Orders

 If beam goes away for more than 2h, or if during the downtime, MCC does operations
 that may change the beam tuning, then do [https://hallcweb.jlab.or/wiki/index.php/Beam_and_Raster_Checkout_for_A1n/d2n_(Short_Procedure)
 beam recovery procedure] (do YAG viewer part only if you are suspicious of
 the harp scan results)
  • NOTE: Shift crew should NOT give authorization to MCC change ion chamber thresholds. If they ask, tell the Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)
  • Check the position locks are on.
  • Do your best to accommodate Hall A's request for a PITA scan promptly. It is OK to stop the present runs and start new ones in 5--10 minutes after they are done.
  • Check that CODA is running normally (had one strange hangup lately)
  • Check online replays shortly after starting runs, & full replay results too.
  • Make sure all the beam specs are followed (e.g. no raster size below 4mm on 3He cell).
  • Always take 120Hz of EDTM
  • Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does). https://logbooks.jlab.org/entry/3783171. If we have to lower the beam current we actually want because the beam is too unstable, prorate the ABUs.
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
!!                                                     !!
!!  ATTN: SHIFT LEADERS                                !!
!!     Please follow the shift summary template on     !!
!!     the laminated SL instructions.                  !!
!!                                                     !!
!!  ATTN: TARGET OPERATORS                             !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit                         !!
!!                                                     !!
!!     For an example see the following entry:         !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • If any magnet trips, snapshot the associated power supply window and post it to hclog before resetting the interlock.
  • Monitor the beam position on target and call MCC immediately if you see large excursions (1 mm level).
  • Ramp rate lock instructions are here and added to checklist[10]

Bluejeans remote connection information

 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920


Useful Links

Target Polarization Angle Conventions

Be VERY clear here about the convention (beam-right/left/up/downstream) vs 0, 90, 180, 270)
for each run period.  More notes and references to log entries are better.

Polarization Calibration Data

ADD LIST OF CALIBRATION RUNS HERE
AND A TABLE OF RELEVANT CALIBRATION CONSTANTS TO USE FOR WHAT RUNS
Cell Calibration
Type
SHMS Angle
/degree
SHMS Momentum
/GeV
Field Direction
/degree
Calibration Results Status Oven Temperature
/°C
Log#
Tommy EPR-NMR 11 7.5 90 5.24 %/mV Done 200 3839939
Tommy EPR-NMR 18 5.6 270 5.04(1) %/mV Done 205 3843301
Tommy EPR-NMR 18 5.6 90 undone 205
Tommy EPR-NMR 11 7.5 270 undone 208
Tommy AFP Loss 14.5 6.4 90 1.25%/Sweep Done 200 3840618, 3840143, 3840618
Tommy EPR Amplifier
on/off NMR Signal
11 7.5 90 Done 200 3839962
Tommy Temperature Test 18 5.6 90 Done 205 3844089
Tommy AFP Loss 18 5.6 270 Done 205 3845255
Tommy Temperature Test 18 5.6 270 Done 205 3845336
Briana EPR-NMR 18 5.6 180 8.60 Done 205 3834036, 3834243
Briana EPR-NMR 18 5.6 270 5.6 Done 205 3835512, 3835587
Austin EPR-NMR 18 5.6 90 4.88 Done 200 3814717
Austin EPR-NMR 11 7.5 180 7.64 Done 210 3817200
Austin EPR-NMR 11 7.5 0 8.13 Done 210 3818761
Austin EPR-NMR 11 7.5 270 4.56 Done 210 3819882
Austin EPR-NMR 11 7.5 270 4.58 Done 210 3822296
Austin EPR-NMR 11 7.5 90 4.91 Done 210 3824907
Austin EPR-NMR 18 0 90 4.58 Done 210 3815895
Austin EPR-NMR 11 0 180 6.91 Done 210 3816724

Target Polarization Database

d2n Run Period

D2n Target Run List: https://docs.google.com/spreadsheets/d/16TrIUXVA-znVNUyEKy9FHaIlcvq-KXzu97k9nJowuE8/edit#gid=0

  • Junhao Chen is maintaining it

A1n Run Period

https://docs.google.com/spreadsheets/d/e/2PACX-1vSoTyDugD8wt3hKmBGZBxwUJTPHmUgyvGVJvM7N6qeVoNWhEXumCwwkPI7Sw9Asae3N4k9h5L0Y7qmw/pubhtml

  • Mingyu Chen will updating the runlist while Murchhana Roy will do the cross check.

Recent polarization measurements

PolHe3 Polarization Measurements






Misc Links